# Microlensing optical depth, event rate, and limits on compact objects in dark matter based on 20 years of OGLE observations of the Small Magellanic Cloud

### v1

@@ Abstract
Some previous studies have suggested that massive and intermediate-mass primordial black holes (PBHs) could comprise a substantial fraction of dark matter in the Universe. Such black holes, if they existed in the Milky Way halo, would give rise to long-duration microlensing events that may potentially last for years. However, earlier searches were not sufficiently sensitive to detect such events. Here, we present the results of searches for long-timescale gravitational microlensing events toward the Small Magellanic Cloud (SMC) using nearly 20 years of photometric observations collected by the Optical Gravitational Lensing Experiment (OGLE) from 2001 to 2020. We found six events, three of which are new discoveries. We use a sample of five events to measure the microlensing optical depth toward the SMC $\tau=(0.32\pm 0.18)\times 10^{-7}$ and the event rate $\Gamma=(1.18\pm 0.57)\times 10^{-7}\,\mathrm{yr}^{-1}\,\mathrm{star}^{-1}$. The properties of the detected events are consistent with lenses originating from known stellar populations within the SMC or in the Milky Way disk. No events with timescales longer than 1 yr were detected, which provides competitive limits on the fraction of massive compact objects, including PBHs, in the Milky Way dark matter halo. Together with the earlier OGLE studies of microlensing events in the direction of the Large Magellanic Cloud, these observations rule out PBHs and other compact objects with masses ranging from $10^{-8}$ to $10^{3}\,M_{\odot}$ as dominant components of dark matter.
Gravitational microlensing (672), Dark matter (353), Milky Way dark matter halo (1049), Small Magellanic Cloud (1468), Primordial black holes (1292), Intermediate-mass black holes (816)
@@ 1 Introduction
The hypothesis that primordial black holes (PBHs) may account for a substantial part of dark matter has gained some popularity in recent years, especially after the discoveries of mergers of unexpectedly massive black holes by gravitational-wave detectors LIGO and Virgo (e.g., Bird et al., 2016; Sasaki et al., 2016; Clesse & García-Bellido, 2017). PBHs are hypothesized to have formed in the very early Universe by the collapse of large density perturbations (e.g., Zel’dovich & Novikov, 1967; Hawking, 1971; Carr & Hawking, 1974; Chapline, 1975). Because the formation of PBHs is thought to have occurred before the Big Bang nucleosynthesis, PBHs are non-baryonic, and their current abundance is not constrained by primordial elemental abundances (e.g., Navas et al., 2024, and references therein). Although PBHs cannot be directly observed, their existence would manifest in a number of phenomena, including (but not limited to) gravitational microlensing of background stars, emission of Hawking radiation, production of gravitational waves, accretion of gas, impact on the dynamics of dwarf galaxies and wide binary star systems, or cosmological structure formation, see the reviews by Green & Kavanagh (2021), Carr et al. (2021) or Carr et al. (2024).
[…]
@@
Some of the strongest constraints on the abundance of PBHs with masses from $10^{-8}\,M_{\odot}$ to $10^{3}\,M_{\odot}$ come from searches for gravitational microlensing events in the long-term photometric observations of stars in the Large Magellanic Cloud (LMC) by the Optical Gravitational Lensing Experiment (OGLE; Mróz et al., 2024b, c, hereafter Paper 1 and Paper 2, respectively). High-cadence OGLE observations of the LMC and the Small Magellanic Cloud (SMC) provide strong limits on the abundance of planetary-mass PBHs (Mróz et al., 2024a, hereafter, Paper 3). These limits supersede the earlier constraints derived from searches for microlensing events toward the LMC by MACHO (Alcock et al., 1997b, 2000), EROS (Tisserand et al., 2007; Blaineau et al., 2022), OGLE-II (Wyrzykowski et al., 2009), and OGLE-III (Wyrzykowski et al., 2011a) experiments by more than an order of magnitude. However, to probe possible dark matter substructures in the Milky Way halo, it is important to search for gravitational microlensing events toward different sky directions.
@@
One possible target is the Andromeda Galaxy (M31). The searches for massive compact halo objects in the direction of M31 by Calchi Novati et al. (2005) and de Jong et al. (2006) yielded contradictory results due to the difficulty in distinguishing self-lensing events (in which stars from M31 are lensing stars in the same galaxy) from those by PBHs, see also Calchi Novati (2010). On the other hand, the high-cadence observations of M31 by the Subaru Hyper Suprime-Cam (HSC) survey (Niikura et al., 2019) provided strong limits on the abundance of moon- to planetary-mass PBHs ($10^{-10}\!-\!10^{-6}\,M_{\odot}$) in dark matter.
@@
The SMC was also a target of microlensing surveys. The MACHO project, which operated from 1992 to 1999, reported the discovery of two microlensing events toward this galaxy (Alcock et al., 1997a, 1999). The EROS-2 survey (1996–2003) detected only one event. Only one event in the SMC was found in the OGLE-II data (1996–2000) by Wyrzykowski et al. (2010), and three events were found in the OGLE-III data (2001–2009) by Wyrzykowski et al. (2011b). These small numbers were consistent with the expected number of self-lensing events (Calchi Novati et al., 2013; Mróz & Poleski, 2018), in which both the microlensing source and lens are located in the same galaxy. No PBHs were required to explain these events.
@@
Previous SMC surveys were not sensitive to very long-duration microlensing events (with timescales of several years) that are expected from massive and intermediate-mass black holes, similar to those detected in large numbers by the LIGO and Virgo observatories. In this study, we combined the data collected during the third and fourth phases of the OGLE project to create nearly 20-year-long, uniform light curves for about 7.4 million stars in the SMC. In addition, 6.2 million stars were observed by OGLE-IV only. This paper presents the results of our search for the very long-duration microlensing events in this data set.
@@
Note. — The coordinates are given for the epoch J2000. $N_{\rm Epochs}$ is the number of $I$-band epochs, separately for OGLE-III (2001–2009) and OGLE-IV (2010–2020) phases of the survey. $N_{\rm Stars}$ is the number of stars (in thousands) detected in the reference image of a given field. This number is not equal to the number of microlensing source stars that enter the optical depth calculations. We separately provide the number of stars observed during both OGLE-III and OGLE-IV (that is, from 2001 to 2020) and during OGLE-IV only (2010–2020). Columns 6 and 7 may not sum to Column 8 due to rounding. (This table is available in its entirety in machine-readable form.)
@@ 2.1 OGLE-IV Data
The photometric data that are analyzed in this paper were collected during the fourth phase of the OGLE survey (OGLE-IV; Udalski et al. 2015) between June 29, 2010 ($\mathrm{JD}=2,455,377$) and March 15, 2020 ($\mathrm{JD}=2,458,924$), when OGLE operations were halted because of the start of the COVID-19 pandemic. The survey uses a dedicated 1.3-m Warsaw Telescope located at Las Campanas Observatory, Chile. The telescope is equipped with a mosaic CCD camera consisting of 32 detectors, each with a resolution of $2048\times 4102$ pixels. With a pixel scale of $0.26$ arcsec pixel-1, the OGLE-IV camera covers the field of view of approximately 1.4 deg2. Further details about the instrument can be found in Udalski et al. (2015).
The OGLE-IV project observes an area of approximately 190 deg2 surrounding the SMC. This region is divided into 135 fields, whose locations are shown in Figure 1. The equatorial coordinates of the center of each field, along with the number of epochs and the number of detected stars, are provided in Table 1. Most of the data were taken in the $I$-band filter, whose transmission curve closely resembles that of the standard Cousins photometric system. The number of $I$-band epochs varies from 151 to 1090, with a median count of 376. In addition, a smaller number of images were taken through the $V$-band filter (from 7 to 321 per field with a median of 22). The cadence of observations differs depending on the specific field. Fields covering the SMC center were observed at a 2–5 day cadence, whereas the outer fields were observed less frequently, typically every 4 to 6 days. The typical seeing ranged from $0.9$ to $2.0^{\prime\prime}$, with a median seeing of $1.4^{\prime\prime}$. Exposure times varied between 150 and 170 s.
We employed the difference image analysis (DIA) technique (Tomaney & Crotts, 1996; Alard & Lupton, 1998), as implemented by Woźniak (2000), to extract photometry. This method provides high-quality measurements, even in the most crowded stellar regions. For every analyzed field, a reference image is constructed by stacking several (from 2 to 10) low-seeing, low-background individual images. Photometry is performed on difference (subtracted) images, which are generated by subtracting the reference image from the incoming frames. Our photometric pipeline detects and flags stars that are identified in the subtracted images. In addition, forced photometry with fixed centroids is calculated for all stars detected in the reference images. In total, the analyzed fields contain about 13.6 million such stars. Please note that the number of stars that are detected on reference images is not equal to the number of microlensing source stars, as discussed in Section 5.
@@ 2.2 Combined OGLE-III/OGLE-IV Data
Because our main scientific goal was to search for long-duration microlensing events, we supplemented the OGLE-IV light curves with data collected during the third phase of the survey (OGLE-III; Udalski 2003), following the methodology detailed in Paper 1. The OGLE-III data were gathered from June 24, 2001 ($\mathrm{JD}=2,452,085$) to May 2, 2009 ($\mathrm{JD}=2,454,954$). This approach allowed us to extend the OGLE-IV light curves to nearly 20 years.
@@
A smaller number of 2.1 million stars located in the central 2.4 deg2 of the SMC were observed during the second phase of the OGLE survey (OGLE-II; Udalski et al. 1997a; Wyrzykowski et al. 2010) from 1997 to 2000. By combining this data set with the light curves from OGLE-III and OGLE-IV, we can extend the full data set to nearly 24 years. However, the OGLE-II camera had a larger pixel size of $0.417^{\prime\prime}$ compared to its successors, which makes it nearly impossible to extract homogeneous light curves using the methodology developed in Paper 1. Therefore, we use the OGLE-II data, if available, only to vet the microlensing event candidates detected in the combined OGLE-III/OGLE-IV data set.
@@ 3 Search for Microlensing Events
In the first step, we searched for objects that underwent at least one significant brightening (called a ‘bump’) in their light curves. Specifically, we required that at least five consecutive data points had to be brighter than $F_{\rm base}+3\sigma_{\rm base}$, where $F_{\rm base}$ is the mean flux and $\sigma_{\rm base}$ is its standard deviation measured outside the magnified section of the light curve. Additionally, we required that the object was detected by the DIA pipeline on at least three subtracted images during the bump, its amplitude was at least 0.1 mag, and it showed no significant variability beyond the main brightening. This initial filtering reduced the number of candidate events from 13.6 million to 25,973.
@@
In addition to the cuts described in Paper 1, we excluded objects located within $10^{\prime}$ of the globular cluster 47 Tuc, as their photometry was significantly affected by extreme crowding. We also decided to discard high proper motion stars, which displayed significant trends in their light curves that were picked up by our algorithm as microlensing candidates. To this end, we cross-matched the list of selected candidates with the Gaia DR3 archive (Gaia Collaboration et al., 2016, 2021b) and removed any objects with proper motions greater than 10 mas yr-1. All these criteria reduced the number of candidate events to 9614.
@@
High proper motion stars 9614
@@
Event R.A. (J2000) Decl. (J2000) OGLE-IV ID $I_{\rm base}$ (mag) $(V-I)_{\rm base}$ (mag) $E(V-I)$ (mag)
@@
Note. — The table provides the mean observed $I$-band brightness and $V-I$ color in the baseline. The reddening $E(V-I)$ toward each event was taken from Skowron et al. (2021).
Out of all analyzed light curves, only five met all selection cuts. These events constitute the statistical sample that we use for the calculation of the microlensing optical depth, event rate (Section 7), and constraints on compact objects in dark matter (Section 9). Events OGLE-SMC-03 and OGLE-SMC-04 were previously discovered by Wyrzykowski et al. (2011b) in the OGLE-III data, while OGLE-SMC-05, OGLE-SMC-06, and OGLE-SMC-07 are new discoveries. The equatorial coordinates of all events, along with their mean $I$-band magnitudes and $(V-I)$ colors in the baseline, are listed in Table 4.
For completeness, we visually inspected the light curves of all 25,973 objects selected in the first step of our criteria. We found only one additional event, OGLE-SMC-02, which was included in the sample of Wyrzykowski et al. (2011b). This event is a known binary-lens event (Dong et al., 2007), whose light curve significantly differs from that of a simple PSPL model. Hence, it did not satisfy all our selection criteria. In addition, we checked the archival OGLE-II data (1997–2000), which were available for all events except OGLE-SMC-03 and OGLE-SMC-04. We also retrieved the archival MACHO light curves (1992–1999) for all events except OGLE-SMC-02 and OGLE-SMC-07. These additional data did not reveal any suspicious variability, which is consistent with the microlensing classification.
@@
OGLE-SMC-03 $2,454,476.12\pm 0.12$ $41.8^{+5.7}_{-5.2}$ $0.169\pm 0.029$ $+1.80\pm 0.23$ $+0.13\pm 0.13$ $20.18\pm 0.21$ $1.54^{+0.31}_{-0.28}$ 1355.4/1371
OGLE-SMC-03 $2,454,476.19\pm 0.11$ $51.5^{+23.8}_{-13.8}$ $0.117\pm 0.052$ $-0.91\pm 0.56$ $+0.28\pm 0.23$ $20.63^{+0.60}_{-0.44}$ $1.01^{+0.45}_{-0.49}$ 1360.8/1371
OGLE-SMC-03 $2,454,476.14\pm 0.09$ $36.8^{+5.3}_{-4.8}$ $-0.166\pm 0.029$ $+2.01\pm 0.27$ $+0.10\pm 0.15$ $20.20\pm 0.21$ $1.51^{+0.30}_{-0.28}$ 1355.1/1371
OGLE-SMC-03 $2,454,476.20\pm 0.14$ $59.3^{+40.9}_{-20.9}$ $-0.111\pm 0.056$ $-1.04\pm 0.60$ $+0.34\pm 0.27$ $20.70^{+0.69}_{-0.49}$ $0.95^{+0.48}_{-0.52}$ 1360.8/1371
@@
OGLE-SMC-07 $2,458,459.45\pm 0.20$ $18.7^{+2.8}_{-2.3}$ $0.377\pm 0.082$ $+1.95\pm 0.33$ $+0.66\pm 0.37$ $18.68\pm 0.33$ $0.99\pm 0.28$ 1380.2/1401
OGLE-SMC-07 $2,458,459.48\pm 0.21$ $19.3^{+2.9}_{-2.3}$ $0.376\pm 0.084$ $+3.20\pm 0.58$ $-0.04\pm 0.20$ $18.69\pm 0.34$ $0.99\pm 0.29$ 1380.0/1401
OGLE-SMC-07 $2,458,459.43\pm 0.19$ $17.8^{+2.7}_{-2.2}$ $-0.376\pm 0.082$ $+2.15\pm 0.37$ $+0.56\pm 0.36$ $18.69\pm 0.33$ $0.99\pm 0.29$ 1380.2/1401
OGLE-SMC-07 $2,458,459.43\pm 0.18$ $17.3^{+2.7}_{-2.2}$ $-0.375\pm 0.082$ $+3.45\pm 0.65$ $-0.30\pm 0.15$ $18.69\pm 0.33$ $0.99\pm 0.29$ 1380.0/1401
Figure 2: $I$-band light curves of detected microlensing events. The solid blue line shows the best-fit microlensing model. The dotted lines for OGLE-SMC-06 and OGLE-SMC-07 present the best-fit PSPL models without microlensing parallax.

### v2

@@ Abstract
Some previous studies have suggested that massive and intermediate-mass primordial black holes (PBHs) could comprise a substantial fraction of dark matter in the Universe. Such black holes, if they existed in the Milky Way halo, would give rise to long-duration microlensing events that may potentially last for years. However, earlier searches were not sufficiently sensitive to detect such events. Here, we present the results of searches for long-timescale gravitational microlensing events toward the Small Magellanic Cloud (SMC) using nearly 20 yr of photometric observations collected by the Optical Gravitational Lensing Experiment (OGLE) from 2001–2020. We found six events, three of which are new discoveries. We use a sample of five events to measure the microlensing optical depth toward the SMC $\tau=(0.32\pm 0.18)\times 10^{-7}$ and the event rate $\Gamma=(1.18\pm 0.57)\times 10^{-7}\,\mathrm{yr}^{-1}\,\mathrm{star}^{-1}$. The properties of the detected events are consistent with lenses originating from known stellar populations within the SMC or in the Milky Way disk. No events with timescales longer than 1 yr were detected, which provides competitive limits on the fraction of massive compact objects, including PBHs, in the Milky Way dark matter halo. Together with the earlier OGLE studies of microlensing events in the direction of the Large Magellanic Cloud, these observations rule out PBHs and other compact objects with masses ranging from $10^{-8}$$10^{3}\,M_{\odot}$ as dominant components of dark matter.
Gravitational microlensing (672); Dark matter (353); Milky Way dark matter halo (1049); Small Magellanic Cloud (1468); Primordial black holes (1292); Intermediate-mass black holes (816)
@@ 1 Introduction
The hypothesis that primordial black holes (PBHs) may account for a substantial part of dark matter has gained some popularity in recent years, especially after the discoveries of mergers of unexpectedly massive black holes by gravitational-wave detectors LIGO and Virgo (e.g., Bird et al., 2016; Sasaki et al., 2016; Clesse & García-Bellido, 2017). PBHs are hypothesized to have formed in the very early Universe by the collapse of large density perturbations (e.g., Zel’dovich & Novikov, 1967; Hawking, 1971; Carr & Hawking, 1974; Chapline, 1975). Because the formation of PBHs is thought to have occurred before the Big Bang nucleosynthesis, PBHs are nonbaryonic, and their current abundance is not constrained by primordial elemental abundances (e.g., Navas et al., 2024, and references therein). Although PBHs cannot be directly observed, their existence would manifest in a number of phenomena, including (but not limited to) gravitational microlensing of background stars, emission of Hawking radiation, production of gravitational waves, accretion of gas, impact on the dynamics of dwarf galaxies and wide binary star systems, or cosmological structure formation (see the reviews by Green & Kavanagh 2021; Carr et al. 2021; Carr et al. 2024).
Despite this rich phenomenology, searches for PBH signatures have not provided compelling arguments for the PBH dark matter hypothesis. Instead, these investigations have led to numerous constraints on the amount of the PBH dark matter in the Universe as a function of the PBH mass. These constraints are often calculated under the assumption of a smooth distribution of dark matter in galactic halos. However, numerical simulations of galaxy formation indicate that dark matter halos may contain numerous substructures, with characteristic sizes depending on the specific dark matter model (e.g., Bechtol et al., 2022). They may be detected, for example, using observations of strongly lensed galaxies. Vegetti et al. (2010, 2012) reported the discovery of such substructures with masses of the order of $10^{8}\!-\!10^{9}\,M_{\odot}$ but it is unclear whether they are associated with lensing galaxies or some unrelated structures along the line of sight (e.g., Li et al., 2017; Despali et al., 2018; Sengül et al., 2022).
@@
Some of the strongest constraints on the abundance of PBHs with masses from $10^{-8}\,M_{\odot}$ to $10^{3}\,M_{\odot}$ come from searches for gravitational microlensing events in the long-term photometric observations of stars in the Magellanic Clouds, as first proposed by Paczyński (1986). The long-term observations of the Large Magellanic Cloud (LMC) by the Optical Gravitational Lensing Experiment (OGLE; Mróz et al., 2024b, c, hereafter Paper 1 and Paper 2, respectively) demonstrate that less than 1% of dark matter may be composed of compact objects with masses from $1.8\times 10^{-4}$ to $6.3\,M_{\odot}$. High-cadence OGLE observations of the LMC and the Small Magellanic Cloud (SMC) provide strong limits on the abundance of planetary-mass PBHs (Mróz et al., 2024a, hereafter, Paper 3). These limits supersede the earlier constraints derived from searches for microlensing events toward the LMC by MACHO (Alcock et al., 1997b, 2000), EROS (Tisserand et al., 2007; Blaineau et al., 2022), OGLE-II (Wyrzykowski et al., 2009), and OGLE-III (Wyrzykowski et al., 2011a) experiments by more than an order of magnitude. However, to probe possible dark matter substructures in the Milky Way halo, it is important to search for gravitational microlensing events toward different sky directions.
@@
One possible target is the Andromeda galaxy (M31). The searches for massive compact halo objects in the direction of M31 by Calchi Novati et al. (2005) and de Jong et al. (2006) yielded contradictory results due to the difficulty in distinguishing self-lensing events (in which stars from M31 are lensing stars in the same galaxy) from those by PBHs (see also Calchi Novati 2010). On the other hand, the high-cadence observations of M31 by the Subaru Hyper Suprime-Cam (HSC) survey (Niikura et al., 2019) provided strong limits on the abundance of moon- to planetary-mass PBHs ($10^{-10}\!-\!10^{-6}\,M_{\odot}$) in dark matter.
@@
The SMC was also a target of microlensing surveys. The MACHO project, which operated from 1992 to 1999, reported the discovery of two microlensing events toward this galaxy (Alcock et al., 1997a, 1999). The EROS-2 survey (1996–2003) detected only one event. Only one event in the SMC was found in the OGLE-II data (1996–2000) by Wyrzykowski et al. (2010), and three events were found in the OGLE-III data (2001–2009) by Wyrzykowski et al. (2011b). These small numbers were consistent with the expected number of self-lensing events (Sahu, 1994; Calchi Novati et al., 2013; Mróz & Poleski, 2018), in which both the microlensing source and lens are located in the same galaxy, and events by nearby lenses in the Milky Way disk (Gould et al., 1994). No PBHs were required to explain these events.
@@
Previous SMC surveys were not sensitive to very long-duration microlensing events (with timescales of several years) that are expected from massive and intermediate-mass black holes, similar to those detected in large numbers by the LIGO and Virgo observatories. In this study, we combined the data collected during the third and fourth phases of the OGLE project to create nearly 20 yr long, uniform light curves for about 7.4 million stars in the SMC. In addition, 6.2 million stars were observed by OGLE-IV only. This paper presents the results of our search for the very long-duration microlensing events in this data set.
@@
Note. — The coordinates are given for the epoch J2000. $N_{\rm Epochs}$ is the number of $I$-band epochs, separately for OGLE-III (2001–2009) and OGLE-IV (2010–2020) phases of the survey. $N_{\rm Stars}$ is the number of stars (in thousands) detected in the reference image of a given field. This number is not equal to the number of microlensing source stars that enter the optical depth calculations. We separately provide the number of stars observed during both OGLE-III and OGLE-IV (that is, from 2001–2020) and during OGLE-IV only (2010–2020). Columns (6) and (7) may not sum to column (8) due to rounding. (This table is available in its entirety in machine-readable form.)
@@ 2.1 OGLE-IV Data
The photometric data that are analyzed in this paper were collected during the fourth phase of the OGLE survey (OGLE-IV; Udalski et al. 2015) between June 29, 2010 ($\mathrm{JD}=2,455,377$) and March 15, 2020 ($\mathrm{JD}=2,458,924$), when OGLE operations were halted because of the start of the COVID-19 pandemic. The survey uses a dedicated 1.3 m Warsaw Telescope located at Las Campanas Observatory, Chile. The telescope is equipped with a mosaic CCD camera consisting of 32 detectors, each with a resolution of $2048\times 4102$ pixels. With a pixel scale of $0.26$ arcsec pixel-1, the OGLE-IV camera covers the field of view of approximately 1.4 deg2. Further details about the instrument can be found in Udalski et al. (2015).
The OGLE-IV project observes an area of approximately 190 deg2 surrounding the SMC. This region is divided into 135 fields, whose locations are shown in Figure 1. The equatorial coordinates of the center of each field, along with the number of epochs and the number of detected stars, are provided in Table 1. Most of the data were taken in the $I$-band filter, whose transmission curve closely resembles that of the standard Cousins photometric system. The number of $I$-band epochs varies from 151 to 1090, with a median count of 376. In addition, a smaller number of images were taken through the $V$-band filter (from 7 to 321 per field with a median of 22). The cadence of observations differs depending on the specific field. Fields covering the SMC center were observed at a 2–5 day cadence, whereas the outer fields were observed less frequently, typically every 4–6 days. The typical seeing ranged from $0.9$ to $2.0^{\prime\prime}$, with a median seeing of $1.4^{\prime\prime}$. Exposure times varied between 150 and 170 s.
We employed the difference image analysis (DIA) technique (Tomaney & Crotts, 1996; Alard & Lupton, 1998), as implemented by Woźniak (2000), to extract photometry. This method provides high-quality measurements, even in the most crowded stellar regions. For every analyzed field, a reference image is constructed by stacking several (from 2–10) low-seeing, low-background individual images. Photometry is performed on difference (subtracted) images, which are generated by subtracting the reference image from the incoming frames. Our photometric pipeline detects and flags stars that are identified in the subtracted images. In addition, forced photometry with fixed centroids is calculated for all stars detected in the reference images. In total, the analyzed fields contain about 13.6 million such stars. Please note that the number of stars that are detected on reference images is not equal to the number of microlensing source stars, as discussed in Section 5.
@@ 2.2 Combined OGLE-III/OGLE-IV Data
Because our main scientific goal was to search for long-duration microlensing events, we supplemented the OGLE-IV light curves with data collected during the third phase of the survey (OGLE-III; Udalski 2003), following the methodology detailed in Paper 1. The OGLE-III data were gathered from June 24, 2001 ($\mathrm{JD}=2,452,085$) to May 2, 2009 ($\mathrm{JD}=2,454,954$). This approach allowed us to extend the OGLE-IV light curves to nearly 20 yr.
@@
A smaller number of 2.1 million stars located in the central 2.4 deg2 of the SMC were observed during the second phase of the OGLE survey (OGLE-II; Udalski et al. 1997a; Wyrzykowski et al. 2010) from 1997–2000. By combining this data set with the light curves from OGLE-III and OGLE-IV, we can extend the full data set to nearly 24,yr. However, the OGLE-II camera had a larger pixel size of $0.417^{\prime\prime}$ compared to its successors, which makes it nearly impossible to extract homogeneous light curves using the methodology developed in Paper 1. Therefore, we use the OGLE-II data, if available, only to vet the microlensing event candidates detected in the combined OGLE-III/OGLE-IV data set.
@@ 3 Search for Microlensing Events
In the first step, we searched for objects that underwent at least one significant brightening (called a “bump”) in their light curves. Specifically, we required that at least five consecutive data points had to be brighter than $F_{\rm base}+3\sigma_{\rm base}$, where $F_{\rm base}$ is the mean flux, and $\sigma_{\rm base}$ is its standard deviation measured outside the magnified section of the light curve. Additionally, we required that the object be detected by the DIA pipeline on at least three subtracted images during the bump, its amplitude be at least 0.1 mag, and it show no significant variability beyond the main brightening. This initial filtering reduced the number of candidate events from 13.6 million to 25,973.
@@
In addition to the cuts described in Paper 1, we excluded objects located within $10^{\prime}$ of the globular cluster 47 Tuc, as their photometry was significantly affected by extreme crowding. We also decided to discard high proper-motion stars, which displayed significant trends in their light curves that were picked up by our algorithm as microlensing candidates. To this end, we cross-matched the list of selected candidates with the Gaia DR3 archive (Gaia Collaboration et al., 2016, 2021b) and removed any objects with proper motions greater than 10 mas yr-1. All of these criteria reduced the number of candidate events to 9614.
@@
High proper-motion stars 9614
@@
Event R.A. (J2000) Decl. (J2000) OGLE-IV ID $I_{\rm base}$ (mag) $(V-I)_{\rm base}$ (mag) E(VI)E(V-I) (mag)
@@
Note. — The table provides the mean observed $I$-band brightness and VIV-I color in the baseline. The reddening E(VI)E(V-I) toward each event was taken from Skowron et al. (2021).
Out of all analyzed light curves, only five met all selection cuts. These events constitute the statistical sample that we use for the calculation of the microlensing optical depth, event rate (Section 7), and constraints on compact objects in dark matter (Section 9). Events OGLE-SMC-03 and OGLE-SMC-04 were previously discovered by Wyrzykowski et al. (2011b) in the OGLE-III data, while OGLE-SMC-05, OGLE-SMC-06, and OGLE-SMC-07 are new discoveries. The equatorial coordinates of all events, along with their mean $I$-band magnitudes and (VI)(V-I) colors in the baseline, are listed in Table 4.
@@
OGLE-SMC-03 $2,454,476.12\pm 0.12$ $41.8^{+5.7}_{-5.2}$ $0.169\pm 0.029$ $1.80\pm 0.23 +$$0.13\pm 0.13$ $20.18\pm 0.21$ $1.54^{+0.31}_{-0.28}$ 1355.4/1371
OGLE-SMC-03 $2,454,476.19\pm 0.11$ $51.5^{+23.8}_{-13.8}$ $0.117\pm 0.052$ $0.91\pm 0.56$ $0.28\pm 0.23$ $20.63^{+0.60}_{-0.44}$ $1.01^{+0.45}_{-0.49}$ 1360.8/1371
OGLE-SMC-03 $2,454,476.14\pm 0.09$ $36.8^{+5.3}_{-4.8}$ $0.166\pm 0.029$ $2.01\pm 0.27 +$$0.10\pm 0.15$ $20.20\pm 0.21$ $1.51^{+0.30}_{-0.28}$ 1355.1/1371
OGLE-SMC-03 $2,454,476.20\pm 0.14$ $59.3^{+40.9}_{-20.9}$ $0.111\pm 0.056$ $1.04\pm 0.60$ $0.34\pm 0.27$ $20.70^{+0.69}_{-0.49}$ $0.95^{+0.48}_{-0.52}$ 1360.8/1371
@@
OGLE-SMC-07 $2,458,459.45\pm 0.20$ $18.7^{+2.8}_{-2.3}$ $0.377\pm 0.082$ $1.95\pm 0.33 +$$0.66\pm 0.37$ $18.68\pm 0.33$ $0.99\pm 0.28$ 1380.2/1401
OGLE-SMC-07 $2,458,459.48\pm 0.21$ $19.3^{+2.9}_{-2.3}$ $0.376\pm 0.084$ $3.20\pm 0.58$ $0.04\pm 0.20$ $18.69\pm 0.34$ $0.99\pm 0.29$ 1380.0/1401
OGLE-SMC-07 $2,458,459.43\pm 0.19$ $17.8^{+2.7}_{-2.2}$ $0.376\pm 0.082$ $2.15\pm 0.37 +$$0.56\pm 0.36$ $18.69\pm 0.33$ $0.99\pm 0.29$ 1380.2/1401
OGLE-SMC-07 $2,458,459.43\pm 0.18$ $17.3^{+2.7}_{-2.2}$ $0.375\pm 0.082$ $3.45\pm 0.65$ $0.30\pm 0.15$ $18.69\pm 0.33$ $0.99\pm 0.29$ 1380.0/1401
Figure 2: II-band light curves of detected microlensing events. The solid blue line shows the best-fit microlensing model. The dotted lines for OGLE-SMC-06 and OGLE-SMC-07 present the best-fit PSPL models without microlensing parallax.
For completeness, we visually inspected the light curves of all 25,973 objects selected in the first step of our criteria. We found only one additional event, OGLE-SMC-02, which was included in the sample of Wyrzykowski et al. (2011b). This event is a known binary-lens event (Dong et al., 2007), whose light curve significantly differs from that of a simple PSPL model. Hence, it did not satisfy all of our selection criteria. In addition, we checked the archival OGLE-II data (1997–2000), which were available for all events except OGLE-SMC-03 and OGLE-SMC-04. We also retrieved the archival MACHO light curves (1992–1999) for all events except OGLE-SMC-02 and OGLE-SMC-07. These additional data did not reveal any suspicious variability, which is consistent with the microlensing classification.